OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (3, 6) and has a slope of -3/4? (A) y= -3/4x+33/4 (B)y=-3/4x-33/4 (C)y=-3/4+23/4 (D)y=-3/4x-23/4

OpenStudy (ankit042):

y-y1 = m(x-x1)

OpenStudy (ankit042):

m = 3/4 (x1,y1) = 3,6

OpenStudy (anonymous):

which one is that? I have no idea how to do this

OpenStudy (ankit042):

put the values in the equation I mentioned to proceed!

OpenStudy (texaschic101):

you can use point slope form, but you do not have to. You can also just use y = mx + b form.

OpenStudy (anonymous):

yeah but im confused on how to do that

OpenStudy (texaschic101):

y = mx + b slope(m) = -3/4 (3,6) x = 3 and y = 6 now sub and solve for b, the y intercept 6 = -3/4(3) + b 6 = -9/4 + b 6 + 9/4 = b 24/4 + 9/4 = b 33/4 = b so your equation is : y = -3/4x + 33/4

OpenStudy (texaschic101):

it is also possible to solve using point slope form

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = -3/4 (3,6) x1= 3 and y1 = 6 now sub y - 6 = -3/4(x - 3) y - 6 = -3/4x + 9/4 y = -3/4x + 9/4 + 6 y = -3/4x + 9/4 + 24/4 y = -3/4x + 33/4

OpenStudy (texaschic101):

whichever way is easier for you to do

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (anonymous):

no... thank you so much!!!

OpenStudy (texaschic101):

no problem :)